# Is the type-D NUT C-metric really “missing” from the most general Plebański-Demiański solution?

### v1

@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12375053, No. 12205243, No. 11675130, by the Sichuan Science and Technology Program under Grant No. 2023NSFSC1347, and by the Doctoral Research Initiation Project of China West Normal University under Grant No. 21E028.
@@
Appendix: ISM construction of 5-dimensional static solutions with our new normalization scheme
@@
In this appendix, we will show that by using our normalization scheme, the 5-dimensional Schwarzschild-Tangherlini solution can be generated from the flat vacuum background via the ISM construction, contrary to the statement made in Ref. PRD73-044004 . Our derivation strictly follows the spirt of the ISM construction, and is different from that had been done in Ref. PTP114-793 where the Killing metric is modified.
V.1 Generating 5-dimensional Schwarzschild-Tangherlini solution
We first take the 5-dimensional flat Minkowski background
@@ V.1 Generating 5-dimensional Schwarzschild-Tangherlini solution
The generated 5-dimensional static 2-soliton metric is very simple:
@@
This trivial example illuminates that the 5-dimensional Schwarzschild-Tangherlini solution is feasibly generated via ISM procedure with our new normalization scheme from the flat Minkowski vacuum seed.
V.2 Generating 5-dimensional static Emparan-Reall black ring
Next, we want to re-derive the static Emparan-Reall black ring PRL88-101101 solution from the 5-dimensional flat Minkowski background written in another form:
@@ V.2 Generating 5-dimensional static Emparan-Reall black ring
we rewrite $\rho^{2}/\mu_{3}=-\nu_{3}$ and make the replacements: $\mu_{3}\to\mu_{3}-\lambda$, $\nu_{3}\to\nu_{3}-\lambda$, then easily obtain the generating matrix:
@@
The generated 5-dimensional 2-soliton metric is still diagonal:
@@
In the coalescing limit $\mu_{3}\mapsto\mu_{2}$ with $k=1-\nu$, the solution reduces to the 5-dimensional Schwarzschild-Tangherlini solution. On the other hand, in the merging limit $\mu_{2}\mapsto\mu_{1}$ with $k=1-\nu$, the solution coalesces to the above 5-dimensional Minkowski background metric.

### v3

@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12375053, No. 12205243, and No. 11675130, by the Sichuan Science and Technology Program under Grant No. 2023NSFSC1347, and by the Doctoral Research Initiation Project of China West Normal University under Grant No. 21E028.
@@
Appendix: ISM construction of five-dimensional static solutions with our new normalization scheme
@@
In this Appendix, we will show that by using our normalization scheme, the five-dimensional Schwarzschild-Tangherlini solution can be generated from the flat vacuum background via the ISM construction, contrary to the statement made in Ref. PRD73-044004 . Our derivation strictly follows the spirit of the ISM construction, and is different from what had been done in Ref. PTP114-793 where the ansatz of the Killing metric is modified.
V.1 Generating five-dimensional Schwarzschild-Tangherlini solution
We first take the five-dimensional flat Minkowski background
@@ V.1 Generating five-dimensional Schwarzschild-Tangherlini solution
The generated five-dimensional static 2-soliton metric is very simple:
@@
This trivial example illuminates that the five-dimensional Schwarzschild-Tangherlini solution is feasibly generated via the ISM procedure with our new normalization scheme from the flat Minkowski vacuum seed.
V.2 Generating five-dimensional static Emparan-Reall black ring
Next, we want to rederive the static Emparan-Reall black ring PRL88-101101 solution from the five-dimensional flat Minkowski background written in another form:
@@ V.2 Generating five-dimensional static Emparan-Reall black ring
we rewrite $\rho^{2}/\mu_{3}=-\nu_{3}$ and make the replacements, $\mu_{3}\to\mu_{3}-\lambda$, $\nu_{3}\to\nu_{3}-\lambda$, then easily obtain the generating matrix:
@@
The generated five-dimensional 2-soliton metric is still diagonal:
@@
In the coalescing limit $\mu_{3}\mapsto\mu_{2}$ with $k=1-\nu$, the solution reduces to the five-dimensional Schwarzschild-Tangherlini solution. On the other hand, in the merging limit $\mu_{2}\mapsto\mu_{1}$ with $k=1-\nu$, the solution coalesces to the above five-dimensional Minkowski background metric.